OpenStudy (anonymous):

Two quarter circles are inside a square so that it forms an ellipse, each end being at opposite vertices. A side of the square is 4. Find the area of the ellipse and how.

OpenStudy (dumbcow):

Area square = 4*4=16 Area quarter circle = 1/4*pi*r^2 = 1/4*16pi = 4pi The difference is area of corner part of square outside ellipse --> 16 - 4pi Area ellipse = 16 - 2(16-4pi) = 8pi -16

OpenStudy (dumbcow):